OpenStudy (onepieceftw):

What is the sum of the arithmetic sequence 3, 9, 15..., if there are 26 terms? Someone explain what I'm doing wrong here?

OpenStudy (onepieceftw):

\[S_n=\frac{ n }{ 2 }(a_1+a_n)\]

OpenStudy (onepieceftw):

That's the equation, right?

OpenStudy (onepieceftw):

I'm getting 377 but that's not one of the choices on the quiz I'm doing.

OpenStudy (onepieceftw):

What else would it be?

OpenStudy (dtan5457):

My bad, aN should be s(26), you need to find s(26) which is d=6 an=3+(25)6=153 now plug that back in

OpenStudy (dtan5457):

n=26 but aN=153

OpenStudy (onepieceftw):

Okay so it's 2028.

OpenStudy (dtan5457):

yes

OpenStudy (onepieceftw):

Alright, thanks, bud!

OpenStudy (dtan5457):

np man